OpenStudy (anonymous):

Anyone please help me on this?? A catapult launches a boulder with an upward velocity of 112 ft/s. The height of the boulder, h, in feet after t seconds is given by the function h= -16t squared + 112t + 30 . How long does it take the boulder to reach its maximum height? What is the boulder’s maximum height? Round to the nearest hundredth, if necessary. A. 7 s; 30 ft B. 3.5 s; 366 ft C. 3.5 s; 618 ft D. 3.5 s; 226 ft

OpenStudy (anonymous):

\[-16t^2+112t+30\] has its maximum at the vertex, which is always at \(-\frac{b}{2a}\) in your case it is \(-\frac{112}{2\times -16}\)

OpenStudy (anonymous):

i get \(t=\frac{7}{2}=3.5\) to find the maximum height, replace \(t\) by 3.5 and see what you get i would use a calculator

OpenStudy (anonymous):

did you understand how i computed the value of \(t\) that gave \(3.5\)?

OpenStudy (anonymous):

I understand how you got t

OpenStudy (anonymous):

so it's B?